# ScaleGNN: Towards Scalable Graph Neural Networks via Adaptive High-order Neighboring Feature Fusion

### v4

@@ V-A3 Implementation Details
We implement all baseline methods according to their provided code. For three small datasets, we set the epoch number to 100 and select dropout rate, weight decay, and learning rate from {0.1, 0.2, 0.3, 0.4, 0.5, 0.6}, {1e-5, 1e-4, 1e-3} and {0.001, 0.005, 0.01, 0.05, 0.1}, respectively. For three large-scale datasets, we set the epoch number and learning rate to 1000 and 0.001 without weight decay. The embedding dimension $d$ is chosen from {32, 64, 128, 256, 512}, and our method uses GCN as the message aggregator.
@@ V-C2 Contrastive Speedup Analysis
To better illustrate the efficiency advantage, we annotate each baseline method with its relative speedup factor compared to ScaleGNNb. It is evident that methods like GBP and GAMLP exhibit substantially higher computational costs, with runtime exceeding 6× that of ScaleGNNb in certain cases. Even RpHGNN, which demonstrates competitive efficiency, remains 2.5× to 3.1× slower than ScaleGNNb on large datasets. These results highlight the capability of ScaleGNN to effectively mitigate the scalability limitations of existing methods. By reducing redundant computations and dynamically adjusting the information flow, our approach enables substantial efficiency improvements without compromising accuracy. The consistent speedup across diverse datasets further underscores the robustness of our design, making it particularly suitable for large-scale graph learning tasks.
@@ V-E Ablation Study (RQ4)
TABLE IV: The comparison of ScaleGNN and its variants on micro-F1 and macro-F1 (Mi-F1 and Ma-F1 for short) metrics.
@@
The results demonstrate that each component is vital for improving ScaleGNN’s performance. w/o Ada variant experiences the largest performance drop across all datasets, with reductions of 1.77% in Micro-F1 and 1.53% in Macro-F1 on ogbn-arxiv, 1.16% and 1.54% on ogbn-products, and 1.34% and 1.42% on ogbn-paper100M, highlighting the importance of adaptive high-order feature fusion for capturing structural dependencies and mitigating information loss. w/o Low variant also shows notable degradation, especially on ogbn-arxiv and ogbn-paper100M, with Micro-F1 dropping by 0.95% and 0.56%, respectively, indicating that low-order enhanced feature aggregation is crucial for maintaining local information. w/o LCS variant sees moderate degradation, with Micro-F1 decreasing over three large-scale datasets, showing the effectiveness of high-order redundant feature masking in reducing redundancy and preventing excessive feature mixing. w/o SC, showing that it still falls short of the full-fledged ScaleGNN model despite being the best of all variants, demonstrating the important role of the LCS-based sparse constraints. It employs an attention-based relevance function, which allows ScaleGNN to adaptively assess and filter neighbors in a differentiable and task-driven manner, significantly improving its generalization and scalability in large-scale graphs.
@@
(c) ogbn-products
(d) ogbn-papers100M
Figure 5: The parameter sensitivity experimental results.

### v5

@@ V-A3 Implementation Details
We implement all baseline methods according to their provided code. Generally, the size of the hidden embedding $d_{f}$ is 128, and the maximum number $K$ of hops is 20. For three small datasets, we set the epoch number to 100 and select dropout rate, weight decay, and learning rate from {0.1, 0.2, 0.3, 0.4, 0.5, 0.6}, {1e-5, 5e-5, 1e-4, 5e-4, 1e-3, 5e-3, 1e-2} and {0.0001, 0.0005, 0.001, 0.005, 0.01}, respectively. For three large-scale datasets, we set the epoch number to 1000 and the learning rate to 0.001 without weight decay. The embedding dimension $d$ is chosen from {32, 64, 128, 256, 512}, and our method uses GCN as the message aggregator. All experiments are conducted on a machine with an AMD EPYC 7402 24-core CPU and NVIDIA GeForce RTX 4090 (24GB Memory) GPU.
@@ V-C2 Contrastive Speedup Analysis
To better illustrate the efficiency advantage, we annotate each baseline method with its relative speedup factor compared to ScaleGNNb. It is evident that methods like GBP and GAMLP exhibit substantially higher computational costs, with runtime exceeding 6× that of ScaleGNNb in certain cases. Even RpHGNN, which demonstrates competitive efficiency, remains 2.5× to 3.1× slower than ScaleGNNb on large datasets. These results highlight the capability of ScaleGNN to effectively mitigate the scalability limitations of existing methods. By reducing redundant computations and dynamically adjusting the information flow, our approach enables substantial efficiency improvements without compromising accuracy. The consistent speedup across diverse datasets further underscores the robustness of our design, making it particularly suitable for large-scale graph representation learning tasks.
@@ V-E Ablation Study (RQ4)
TABLE IV: The comparison of ScaleGNN and its variants in terms of Mi-F1 and Ma-F1 on three large-scale datasets.
@@
The results demonstrate that each component is vital for improving ScaleGNN’s performance. w/o Ada variant experiences the largest performance drop across all datasets, with reductions of 3.07% in Micro-F1 and 2.17% in Macro-F1 on ogbn-arxiv, 2.07% and 1.14% on ogbn-products, and 1.24% and 1.44% on ogbn-paper100M, highlighting the importance of adaptive high-order feature fusion for capturing structural dependencies and mitigating information loss. w/o Low variant also shows notable degradation, especially on ogbn-arxiv and ogbn-paper100M, with Micro-F1 dropping by 1.33% and 0.72%, respectively, indicating that low-order enhanced feature aggregation is crucial for maintaining local information. w/o LCS variant sees moderate degradation, with Micro-F1 decreasing over three large-scale datasets, showing the effectiveness of high-order redundant feature masking in reducing redundancy and preventing excessive feature mixing. w/o SC shows that it still falls short of the full-fledged ScaleGNN model despite being the best of all variants, demonstrating the important role of the LCS-based sparse constraints. It employs an attention-based relevance function, which allows ScaleGNN to adaptively assess and filter neighbors in a differentiable and task-driven manner, significantly improving its generalization and scalability in large-scale graphs.
@@
Figure 5: The effect of balanced hyperparameter $\beta$.
(a) ogbn-products
(b) ogbn-papers100M
Figure 6: The effect of loss fuction weights $\lambda_{1}$ and $\lambda_{2}$.